OpenStudy (anonymous):

A line has a vertical change per unit of horizontal change of negative five-fourths and passes through point two comma two. Which is another point on the line?

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

@ParthKohli @Prince1342

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

@RadEn

Parth (parthkohli):

Hi! Have you heard the term "slope" in the context of a line?

Parth (parthkohli):

Well, never mind. You won't need it.

Parth (parthkohli):

It is given that (2,2) is a point on the line. If x increases by 1, then y increases by -4/5. That is what this question means.

OpenStudy (anonymous):

?? I don't get it. How do I find the answer?

Parth (parthkohli):

sorry, let me explain it now. I was on another question.

OpenStudy (anonymous):

Ok

Parth (parthkohli):

"A line has a vertical change per unit of horizontal change of negative five-fourths" Let's focus on what this means for a moment.

Parth (parthkohli):

so this statement means this: if there is a horizontal change of 1 on the line then there is a vertical change of -5/4 on the same line

Parth (parthkohli):

horizontal means the x axis, right?

OpenStudy (anonymous):

Hmm

OpenStudy (anonymous):

Yeah, but how do you find the answer?

Parth (parthkohli):

I'm sorry, I must go now.

OpenStudy (anonymous):

Ok...

OpenStudy (anonymous):

@sean_davis0821 Do you know?

OpenStudy (anonymous):

Sorry I'm really bad with those kind of problems @AnswerMyQuestions

OpenStudy (anonymous):

Oh, Okay.

OpenStudy (anonymous):

Sorry again

OpenStudy (anonymous):

It's Ok.

OpenStudy (phi):

this is pretty obtuse language : "A line has a vertical change per unit of horizontal change of negative five-fourths " what it is saying is change in y is -5/4 and change in x is 1 change in y divided by change in x is SLOPE of the line -5/4 divided by 1 is -5/4 (anything divided by 1 is itself) that means slope is -5/4 the equation of a line in point slope form is y - y0 = m(x - x0) where m is slope and (x0,y0) is a point on the line

OpenStudy (phi):

they tell you the line " passes through point two comma two. " meaning (2,2) is a point on the line. you can write down the equation of this line, using (2,2) = (x0,y0) and m= -5/4 put those numbers into the formula y - y0 = m(x - x0) can you do that ?

OpenStudy (anonymous):

I don't know.

OpenStudy (phi):

if you are given a formula y = A*x and are told A=3. you should erase the A, and put 3 in its place to get y = 3*x

OpenStudy (phi):

here you start with a formula y - y0 = m(x - x0) and are told m= -5/4 replace the m with -5/4 in the formula. can you do that ?

OpenStudy (anonymous):

y - y0 = -5/4(x - x0)

OpenStudy (phi):

yes. also (x0,y0) is (2,2) that means x0=2 and y0=2. can you replace x0 and y0 in the formula y - y0 = -5/4(x - x0)

OpenStudy (anonymous):

y - 2 = -5/4(x - 2)

OpenStudy (phi):

yes, and that is the equation of a line in point slope form. you can use it to find other points on the line. Are you given a list of points as possible answers or do you just have to find another point on the line ?

OpenStudy (anonymous):

I have choices.

OpenStudy (phi):

in that case we can replace x and y with the numbers in a choice, and see if it works in the equation

OpenStudy (anonymous):

The choices are (-3,6), (-10,8), (8,-10), and (6,-3)

OpenStudy (anonymous):

@phi What next?

OpenStudy (phi):

try the first choice (-3,6). (-3,6) is short for x= -3 and y=6 we use those numbers for x and y in your formula y - 2 = -5/4(x - 2) replace y with 6 6 - 2 = -5/4 (x - 2) replace x with -3 6 -2 = -5/4 ( -3 - 2) can you simplify both sides ? on the left you get 6-2 = ?

OpenStudy (anonymous):

6 - 2 = 4

OpenStudy (phi):

ok so we have 4 = -5/4 ( -3 - 2) can you simplify the right side ? -3 - 2 is - 5 4= -5/4 * -5 now finish the right side. what do we get ?

OpenStudy (anonymous):

6 1/4

OpenStudy (anonymous):

@phi

OpenStudy (phi):

and 4 does not equal 6 ¼ choice A is NOT the answer.

OpenStudy (phi):

to answer this question we have to be good at arithmetic. test the last the choice. can you do that ?

OpenStudy (anonymous):

I'll try.

OpenStudy (phi):

y - 2 = -5/4(x - 2) and (6,-3)